# Think-on-Graph 2.0: Deep and Faithful Large Language Model Reasoning with Knowledge-guided Retrieval Augmented Generation

### v4

@@ Abstract
Retrieval-augmented generation (RAG) has improved large language models (LLMs) by using knowledge retrieval to overcome knowledge deficiencies. However, current RAG methods often fall short of ensuring the depth and completeness of retrieved information, which is necessary for complex reasoning tasks. In this work, we introduce Think-on-Graph 2.0 (ToG-2), a hybrid RAG framework that iteratively retrieves information from both unstructured and structured knowledge sources in a tight-coupling manner. Specifically, ToG-2 leverages knowledge graphs (KGs) to link documents via entities, facilitating deep and knowledge-guided context retrieval. Simultaneously, it utilizes documents as entity contexts to achieve precise and efficient graph retrieval. ToG-2 alternates between graph retrieval and context retrieval to search for in-depth clues relevant to the question, enabling LLMs to generate answers. We conduct a series of well-designed experiments to highlight the following advantages of ToG-2: 1) ToG-2 tightly couples the processes of context retrieval and graph retrieval, deepening context retrieval via the KG while enabling reliable graph retrieval based on contexts; 2) it achieves deep and faithful reasoning in LLMs through an iterative knowledge retrieval process of collaboration between contexts and the KG; and 3) ToG-2 is training-free and plug-and-play compatible with various LLMs. Extensive experiments demonstrate that ToG-2 achieves overall state-of-the-art (SOTA) performance on 6 out of 7 knowledge-intensive datasets with GPT-3.5, and can elevate the performance of smaller models (e.g., LLAMA-2-13B) to the level of GPT-3.5’s direct reasoning.
@@ 1 Introduction
Figure 1: The illustration of differences among (a) text-based RAG, (b) KG-based RAG, (c) KG+Text Hybrid RAG, and (d) our proposed KG×$\times$×Text RAG framework.
Retrieval-augmented generation (RAG) has significantly enhanced the capabilities of large language models (LLMs) by retrieving relevant knowledge from external sources, as a promising solution to address the knowledge deficiencies and hallucination issues(Zhao et al., 2024). Despite this potential, and researchers’ attempts to incorporate various complex additional processes into RAG (Yu et al., 2023; Edge et al., 2024; Li et al., 2024), LLMs still struggle to maintain human-like reasoning trajectories when handling complex tasks, which require continuous effort and motivation to integrate fragmented information and the structural relationships between them.
@@ 3.2.1 Knowledge-guided Graph Search
Entity Discovery: Given a topic entities $e_{j}^{i}$ in $\mathcal{E}_{topic}^{i}$ and its corresponding selected relations $r_{j,m}^{i}$ in $\mathcal{R}^{i}$, function
@@ 3.3 Reasoning with Hybrid Knowledge
The details of the prompts are shown in Appendix E Table 13 and 14.
@@ 4.2 Baselines
We compare ToG-2 with both widely-used baselines and state-of-the-art methods to provide a more comprehensive overview: 1) LLM-only methods without external knowledge, including Direct Reasoning, Chain-of-Thought ((Wei et al., 2022b)), Self-Consistency ((Wang et al., 2023a)); 2) Vanilla RAG, indicating the text-based RAG method that directly retrieves from entity documents and answers the question; 3) KG-based RAG method: Think-on-Graph ((Sun et al., 2024)), a KG-based RAG method that searches useful KG triples for reasoning; 4) Chain-of-Knowledge ((Li et al., 2024)), a hybrid RAG method retrieving knowledge from Wikipedia, Wikidata and Wikitable; 5) GraphRAG ((Edge et al., 2024)), a hybrid RAG method that first builds a KG from documents and enables KG-enhanced text retrieval. Notably, we only evaluate GraphRAG on ToG-FinQA due to its unaffordable computation cost during the indexing process on a large corpus. For a fair comparison, all baselines are used with GPT-3.5-turbo and evaluated under an unsupervised setting.
@@ 4.3 Implementation Details
In this study, we mainly use GPT-3.5-turbo as the backbone LLM for a fair comparison to other baselines. To evaluate the effect of backbone LLMs on ToG-2’s performance, we also conduct experiments with GPT-4o, Llama3-8B and Qwen2-7B. Consistent with the ToG settings, we set the temperature parameter to 0.4 during exploration and 0 during reasoning. The width $W=3$ and the maximum number of iterations (depth) is $3$. For context retrieval, we utilize the pre-trained BGE-embedding model without any fine-tuning. In entity prune, we maintain 10 ($K=10$) top-scored sentences for calculating entity scores. In the reasoning stage, we employ a 2-shot demonstration for most of the datasets. Following CoK (Li et al., 2024), we employ 3-shot and 6-shot demonstration for FEVER and make the LLM execute a self-consistency reasoning step first.
@@ 4.6 Runtime Analysis
In the above-mentioned experiments, we use Equation 3 in relation pruning for efficiency. Under this setting, the reasoning process of ToG-2 needs at most $12D+$ 1 calls to the LLM. When using Equation 2 in relation pruning, ToG-2 needs at most $WD+D+1$ LLM calls and can achieve slightly higher performances on several datasets as shown in Appendix B.2 Table 6. In Table 6, we also report the results of an additional ablation study. By contrast, ToG needs at most $12WD+D+$ 1 LLM calls. We compare the specific runtime between ToG and ToG-2 on different datasets in Appendix B.3 Table 7. Even though ToG-$2$ integrates additional knowledge through entity context, ToG-$2$ still manages to reduce the entity pruning phase to 68.7% of ToG’s runtime on average. This is because ToG relies on LLM-based entity pruning, while ToG-$2$ leverages DRMs, which can handle more candidates and compute relevance scores faster.
@@ Appendix A Pseudocode
       // Knowledge-guided Graph Search
@@ B.1 Manual Analysis (continue)
B.2 Supplementary Ablation Studies
@@
Table 6: Influence of each module in ToG-$2$ on the final performance. / denote without. $Clue$: the retrieval feedback. $TP$: topic prune. $RC$: relation prune combination with Eq.3.
B.3 Runtime Analysis Results
Does ToG-2 improve graph exploring efficiency compared to ToG? Therefore, we statistically compared the runtime of ToG and ToG-$2$ across different stages under the same settings. The results in Table 7 show a substantial improvement in ToG-$2$’s relation pruning time, reducing it to just 45% of ToG’s. This gain comes from the relation pruning combination strategy, which cuts down the number of LLM calls from W times to 1 time each iteration. Even though ToG-$2$ integrates additional knowledge through entity context, ToG-$2$ still manages to reduce the entity pruning phase to 68.7% of ToG’s runtime on average across both datasets. This is because ToG relies on LLM-based entity pruning, while ToG-$2$ leverages DRMs, which can handle more candidates and compute relevance scores faster.
@@ B.3 Runtime Analysis Results
Table 7: Running time comparison of ToG-2 and ToG on FEVER and AdvHotpotQA datasets
B.4 Impact of SC setting
@@ Appendix E Prompts
Table 11: Topic Prune CMB Prompt
@@
Table 12: Topic Prune CMB Prompt (Continue)

### v6

@@ Abstract
Retrieval-augmented generation (RAG) has improved large language models (LLMs) by using knowledge retrieval to overcome knowledge deficiencies. However, current RAG methods often fall short of ensuring the depth and completeness of retrieved information, which is necessary for complex reasoning tasks. In this work, we introduce Think-on-Graph 2.0 (ToG-2), a hybrid RAG framework that iteratively retrieves information from both unstructured and structured knowledge sources in a tight-coupling manner. Specifically, ToG-2 leverages knowledge graphs (KGs) to link documents via entities, facilitating deep and knowledge-guided context retrieval. Simultaneously, it utilizes documents as entity contexts to achieve precise and efficient graph retrieval. ToG-2 alternates between graph retrieval and context retrieval to search for in-depth clues relevant to the question, enabling LLMs to generate answers. We conduct a series of well-designed experiments to highlight the following advantages of ToG-2: 1) ToG-2 tightly couples the processes of context retrieval and graph retrieval, deepening context retrieval via the KG while enabling reliable graph retrieval based on contexts; 2) it achieves deep and faithful reasoning in LLMs through an iterative knowledge retrieval process of collaboration between contexts and the KG; and 3) ToG-2 is training-free and plug-and-play compatible with various LLMs. Extensive experiments demonstrate that ToG-2 achieves overall state-of-the-art (SOTA) performance on 6 out of 7 knowledge-intensive datasets with GPT-3.5, and can elevate the performance of smaller models (e.g., LLAMA-2-13B) to the level of GPT-3.5’s direct reasoning. The source code is available on https://github.com/IDEA-FinAI/ToG-2.
@@ 1 Introduction
Figure 1: The illustration of differences among (a) text-based RAG, (b) KG-based RAG, (c) KG+Text (Loose-coupling) Hybrid RAG, and (d) our proposed KG×$\times$×Text (Tight-coupling) RAG framework.
Retrieval-augmented generation (RAG) has significantly enhanced the capabilities of large language models (LLMs) by retrieving relevant knowledge from external sources, as a promising solution to address the knowledge deficiencies and hallucination issues (Zhao et al., 2024). Despite this potential, and researchers’ attempts to incorporate various complex additional processes into RAG (Yu et al., 2023; Edge et al., 2024; Li et al., 2024), LLMs still struggle to maintain human-like reasoning trajectories (kahneman2011thinking) when handling complex tasks, which require continuous effort and motivation to integrate fragmented information and the structural relationships between them.
@@ 3.2.1 Knowledge-guided Graph Search
Entity Discovery: Given a topic entity $e_{j}^{i}$ in $\mathcal{E}_{topic}^{i}$ and its corresponding selected relations $r_{j,m}^{i}$ in $\mathcal{R}^{i}$, function
@@ 3.3 Reasoning with Hybrid Knowledge
The details of the prompts are shown in Appendix E Table 18 and 19.
@@ 4.2 Baselines
We compare ToG-2 with both widely-used baselines and state-of-the-art methods to provide a more comprehensive overview: 1) LLM-only methods without external knowledge, including Direct Reasoning, Chain-of-Thought (Wei et al., 2022b), Self-Consistency (Wang et al., 2023a); 2) Vanilla RAG, indicating the text-based RAG method that directly retrieves from entity documents and answers the question; 3) KG-based RAG method: Think-on-Graph (Sun et al., 2024), a KG-based RAG method that searches useful KG triples for reasoning; 4) Chain-of-Knowledge (Li et al., 2024), a hybrid RAG method retrieving knowledge from Wikipedia, Wikidata and Wikitable; 5) GraphRAG (Edge et al., 2024), a hybrid RAG method that first builds a KG from documents and enables KG-enhanced text retrieval. Notably, we only evaluate GraphRAG on ToG-FinQA due to its unaffordable computation cost during the indexing process on a large corpus. For a fair comparison, all baselines are used with GPT-3.5-turbo and evaluated under an unsupervised setting.
@@ 4.3 Implementation Details
We mainly use GPT-3.5-turbo as the backbone LLM for a fair comparison to other baselines. To evaluate the effect of backbone LLMs on ToG-2’s performance, we conduct experiments with GPT-4o, Llama3-8B and Qwen2-7B. Consistent with the ToG settings, we set the temperature parameter to 0 for all generations. The width $W=3$ and the maximum number of iterations (depth) is $3$. During relation prune, we set a threshold of 0.2 to filter relations with low relevance scores and then select the top $W$. For context retrieval, we utilize the BGE-embedding model without any fine-tuning. In entity prune, we maintain 10 ($K=10$) top-scored sentences for calculating entity scores. In the reasoning stage, we employ a 2-shot demonstration for most of the datasets. Following CoK (Li et al., 2024), we employ 3-shot and 6-shot demonstration for FEVER and make the LLM execute a self-consistency reasoning step first.
@@ 4.6 Runtime Analysis
In the above-mentioned experiments, we use Equation 3 in Relation Prune for efficiency. Under this setting, the reasoning process of ToG-2 needs at most $12D+(D-1)+$ 1 calls to the LLM. When using Equation 2 in relation pruning, ToG-2 needs at most $WD+D+1$ LLM calls and can achieve slightly higher performances on several datasets as shown in Appendix B.3 Table 7. In Table 7, we also report the results of an additional ablation study. By contrast, ToG needs at most $12WD+D+$ 1 LLM calls. We compare the specific runtime between ToG and ToG-2 on different datasets in Appendix B.4 Table 8. Even though ToG-$2$ integrates additional knowledge through entity context, ToG-$2$ still manages to reduce the entity pruning phase to 68.7% of ToG’s runtime on average. This is because ToG relies on LLM-based entity pruning, while ToG-$2$ leverages DRMs, which can handle more candidates and compute relevance scores faster.
@@ Appendix A Pseudocode
       // Context Enhanced Graph Retrieval
@@ B.1 Manual Analysis (continue)
B.2 Supplementary Analysis for The Main Results
As we observed a performance discrepancy between ToG and ToG-$2$ on the Creak dataset in Table 1, we aim to provide a more detailed analysis on this particular dataset. This analysis will offer insights into the underlying factors contributing to the observed discrepancy. Furthermore, we will showcase and discuss the implications of these findings for understanding the strengths and limitations of both methods in the context of the Creak dataset. Specifically, we manually analyzed the reasoning process for 50 examples. Table 6 provides an overview of the distribution of answer types for ToG-$2$, following the format of Table 4 in our paper.
Answer Type ToG-$2$
Triple-enhanced Answer 42%
Doc-enhanced Answer 24%
Both-enhanced Answer 34%
Table 6: Distribution of answer types for ToG-$2$ on the Creak dataset.
Due to the relatively simple nature of the questions in the Creak dataset, results can often be inferred directly using triples. As a result, ToG already performs well on this dataset, and the richer entity document information utilized by ToG-$2$ does not provide significant additional benefits. As shown in the table, ToG-$2$ produces a higher proportion of Triple-enhanced Answers on Creak compared to more complex datasets like HotpotQA.
We conducted a detailed case study of the behavior of ToG and ToG-$2$, highlighting three representative cases that elucidate the primary reasons for the observed performance differences. These bad cases reveal distinct patterns in how each model approaches reasoning.
  Case 1: One can watch Bugs Bunny on the television.
ToG-2 Response: {Yes}, Bugs Bunny can be watched on television. The references mention the history and success of the Bugs Bunny series, indicating its presence on various channels and streaming services over the years. The knowledge triplets also confirm Bugs Bunny’s association with iconic works like Looney Tunes and notable creators like Tex Avery and Mel Blanc.
ToG Response: {Yes}, the claim is {True}. Based on the knowledge triplets, Bugs Bunny is a beloved character from iconic works such as Looney Tunes, which was created by Warner Bros, and has been featured extensively in television broadcasts since its creation.
Label: True
Error Reason: The model failed to produce the expected output format {True}. The LLM did not strictly follow the instructions, which is a low-probability event but accounts for a large proportion of errors.
Case 2: Quark served the man a drink on Star Trek: Deep Space Nine.
ToG-2 Response: {No}, the claim is {False}. The claim that “Quark served the man a drink on Star Trek: Deep Space Nine” cannot be evaluated based on the given information. The references provide information about the character Quark is a Ferengi bartender on Deep Space Nine, but there is no specific mention of him serving a drink to a man.
ToG Response: {Yes}, the claim is {True}. The knowledge triplets confirm Quark’s association with Star Trek: Deep Space Nine and his role as a bartender, serving drinks to patrons in the series.
Label: True
Error Reason: The model’s failure stems from two main issues: overcautiousness and semantic ambiguity. Despite identifying the golden sentence, “Quark was the bartender on Deep Space Nine,” ToG-$2$’s excessive caution prevented it from making implicit inferences without explicit evidence. Additionally, the phrase “the man” implies a specific individual without sufficient context, leading to a misinterpretation.
Case 3: Parkinson’s disease causes various hallucinations in people.
ToG-2 Response: {Yes}, the claim is {True}. The retrieved references mention that Parkinson’s disease psychosis (PDP) is present in around 20 percent of cases and comprises hallucinations, illusions, and delusions. This confirms that Parkinson’s disease can indeed cause various hallucinations in people.
ToG Response: {No}, the claim is {False}. The retrieved references do not provide any evidence linking Parkinson’s disease to causing various hallucinations in people. As no supporting information was found, it is concluded that this condition does not lead to hallucinations.
Label: False
Error Reason: Although ToG-$2$’s response appears reasonable from a medical perspective, the golden label is “False.” This misalignment suggests that ToG-$2$’s detailed document analysis can sometimes lead to conclusions that diverge from the dataset’s annotations, particularly when the dataset lacks nuanced inferences.
Conclusion: ToG performs well in scenarios that demand rapid, triplet-based reasoning, while ToG-$2$ demonstrates its strength in cases requiring detailed contextual reasoning. However, ToG-$2$’s overcautious behavior and occasional divergence from the ground truth reveal its tendency to prevent hallucinations at the cost of misalignment with less explicit labels. Additionally, the quality issues within some questions in the Creak dataset also limit the upper bound of model performance.
B.3 Supplementary Ablation Studies
@@
Table 7: Influence of each module in ToG-$2$ on the final performance. / denote without. $Clue$: the retrieval feedback. $TP$: topic prune. /$RC$: replace relation prune Eq.3 with Eq.2.
B.4 Runtime Analysis Results
Does ToG-2 improve graph exploring efficiency compared to ToG? Therefore, we first statistically compared the runtime of ToG and ToG-$2$ across different stages under the same settings. Specifically, we analyze the average number of iterations in answering each discrete question, along with the average runtime for each iteration.
The results in Table 8 show a substantial improvement in ToG-$2$’s relation pruning time, reducing it to just 45% of ToG’s. This gain comes from the relation pruning combination strategy, which cuts down the number of LLM calls from W times to 1 time each iteration. Even though ToG-$2$ integrates additional knowledge through entity context, ToG-$2$ still manages to reduce the entity pruning phase to 68.7% of ToG’s runtime on average across both datasets. This is because ToG relies on LLM-based entity pruning, while ToG-$2$ leverages DRMs, which can handle more candidates and compute relevance scores faster.
@@ B.4 Runtime Analysis Results
Table 8: Running time comparison of ToG-2 and ToG on FEVER and AdvHotpotQA datasets.
To better illustrate the efficiency of our method compared to the baselines, we analyze the average total runtime and API call counts for each method on HotpotQA. The results are summarized in Table 9 below.
Method Avg Total Time (s) API Calls
ToG-$2$ 27.3 5.4
NaiveRAG 10.2 1
ToG 69.3 16.3
CoK 30.1 11
Table 9: Comparison of API Call Counts and Runtime Across Methods.
ToG: For each question, ToG requires up to $12WD+D+$ 1 API calls, where $W$ is the maximum width (number of knowledge triplets explored in parallel) and $D$ is the maximum iteration depth. This structure enables comprehensive multi-hop reasoning and knowledge exploration, but results in a relatively high number of API calls per question.
CoK: To ensure fairness in comparison, we restricted the knowledge domain to factual data from WikiData, Wikipedia, and DPR. We also adopted the update_rationales_at_once setting from the original implementation. CoK has a fixed API call count of 11 per question: 1 domain selection call, 1 chain-of-thought sub-question generation call, $2\times 3$ knowledge retrieval calls, 2 rationale editing calls, and 1 final answer generation call.
NaiveRAG: To simulate a retrieval corpus, we use all entity documents encountered during ToG-$2$ iterations. NaiveRAG uses a two-stage retrieval system that combines BM25 with BGE-Reranker, resulting in the fewest API calls among the compared methods.
ToG-$2$: For each question, ToG-$2$ requires up to $12D+(D-1)+$ 1 API calls, where:
  • $2D$: Calls for Relation Pruning (RP) and Reasoning during each iteration.
  • $D-1$: Query re-writing between iterations.
  • $1$: Topic Prune (TP).
Compared to baselines, ToG-$2$ achieves a good balance between cost-efficiency and performance. It reduces the total API calls and runtime while maintaining robust reasoning capabilities.
B.5 Impact of Threshold Setting in Relation Prune
We conducted threshold experiments in Relation Prune (RP) to analyze the effect of different threshold scores. Specifically, we compared the effects of three thresholds (0.2, 0.5, and 0.8) on the final results in the HotpotQA dataset. The results are summarized in Table 10.
Threshold Results (%) Avg. Retained Score
0.2 42.9 0.65
0.5 43.2 0.68
0.8 40.0 0.92
Table 10: Effects of different thresholds on the final results in HotpotQA. The Avg. retained score represents the average score of all relations that pass through the threshold filtering.
When the threshold is set between 0.2 and 0.5, its impact on the final result is minimal. However, when the threshold increases to 0.8, the results weaken significantly. We attribute this to two main reasons:
  1. 1.
    During the relation pruning stage, the relations passing the threshold are further filtered by selecting the top relations based on the width $W$. Therefore, even if the threshold is set very low, relations with extremely low scores will ultimately fail to be included in the Top $W$. This explains why the Avg. retained score is very close when the threshold is 0.2 or 0.5. However, setting a reasonable threshold is still necessary to prevent the inclusion of entirely irrelevant relations in cases where very few relations are found and are scored low (e.g., as low as 0.1).
  2. 2.
    Relations with moderate relevance scores (e.g., 0.3–0.7) can potentially lead to useful information. A lower threshold allows for more tolerance for such potential. When the threshold is raised to 0.8, it overly amplifies the impact of the LLM’s judgment during relation pruning on the entire process, leading to the loss of potentially valuable relations.
B.6 Impact of SC Setting
@@
B.7 Impact of Graph Completeness
To investigate the impact of graph quality on the reasoning performance of ToG-$2$, we designed an experiment to quantify the specific effect of Knowledge Graph (KG) quality on the model’s performance. Specifically, we conducted experiments on a subset of randomly sampled HotpotQA questions (100 in total) using simulated incomplete KGs. The incompleteness was introduced by randomly discarding relations and entities discovered during the Relation Discovery (Equation 1) and Entity Discovery (Equation 4) stages. Specifically, we tested scenarios where 30%, 50%, and 80% of the discovered entities and relations were retained, while the rest were randomly removed. This systematic evaluation provided insights into the model’s adaptability to varying levels of KG completeness. The results are shown in Table 11.
KG Completeness (%) Exploration Setting EM (%)
100 Default 43
80 Default 41
50 Default 35
30 Default 23
30 Adjusted 29
Table 11: Impact of KG Completeness on ToG-2.0 Performance with Default ($W=3$, $D=3$) and Adjusted ($W=8$, $D=2$) Exploration Strategies
At 80% completeness, the model demonstrated robust performance with minimal impact compared to the fully complete KG, indicating its strength at high completeness levels. With 50% completeness, moderate performance degradation was observed, suggesting that ToG-2.0 remains resilient to moderate KG incompleteness. However, at 30% completeness, performance dropped significantly due to sparse connections in the graph, highlighting the challenges posed by severe KG incompleteness.
To mitigate this, adjustments were made to the exploration strategy: By increasing the exploration width ($W$) to 8 and reducing the exploration depth ($D$) to 2, the model was able to broaden its search scope, compensating for the missing information. This adjustment led to a notable recovery in performance, bringing it closer to an acceptable level despite the significant KG incompleteness. These findings emphasize the model’s adaptability and provide insights into potential strategies for handling incomplete KGs.
B.8 Significance Test
Statistical significance is a key factor in evaluating model performance. To assess this, we conducted pairwise t-tests on select datasets where we had access to both ToG-2 and baseline results. However, many of the baseline results reported in Table 1 were directly cited from their respective original papers or other related works. As a result, we do not have access to the detailed outputs of these baseline models for individual questions, preventing us from performing significance tests on those baselines at this time.
For the WebQSP, Zero-Shot-RE, QALD-10-en, and Creak datasets, we computed pairwise t-tests comparing ToG-2 with CoK and ToG as in Table 12.
(ToG-2 vs. ) Method WebQSP Zero-Shot-RE QALD-10-en Creak
CoK $p<0.05$ $p<0.01$ $p<0.05$ $p<0.05$
ToG $p<0.05$ $p<0.05$ $p<0.05$ $p>0.9$
Table 12: ToG-2 significantly outperforms CoK and ToG on most datasets ($p<0.05$), except on Creak where the difference with ToG is not significant ($p>0.9$), indicating similar performance on this dataset.
@@ Appendix E Prompts
Table 16: Relation Prune CMB Prompt
@@
Table 17: Relation Prune CMB Prompt (Continue)